# Large Language Models on Fine-grained Emotion Detection Dataset with Data Augmentation and Transfer Learning

### 10 Mar 2024 v1

@@ Abstract
1 Problem Description
Recently Large Language Models have gain enormous attention in the field of Natural Language Processing. The task of detecting subtle emotions in text is a complex issue in Natural Language Processing with numerous practical uses. A major hurdle is the lack of sufficiently large and annotated datasets for emotion detection. Demszky Demszky et al. (2020) created the GoEmotions dataset, which is the largest manually annotated dataset containing 58k English Reddit comments, labeled for 27 emotion categories and Neutral. The researchers also fine-tuned a BERT model on GoEmotions as a strong baseline in transfer learning across domains and taxonomies to show the potentials of this dataset.
In the initial experiments, our team dived deep into the original paper and reproduced the results of both fine-tuning BERT model on GoEmotions dataset and applying transfer learning technique using fine-tuned BERT to other emotion classification dataset with different taxonomies. We were able to achieve 0.49 macro-average F1 score on the 28-label classification task, which was slightly improved over the best result 0.46 reported in the original paper Demszky et al. (2020). However, we believe that there is still much room for improvement in term of classification performance on GoEmotions dataset. Therefore, our main objective in this study is to utilize different methods to enhance the classification performance on the fine-grained emotion detection dataset, GoEmotions, as much as possible.
@@ 2 Related Work
Our study builds upon the foundation laid by GoEmotions Demszky et al. (2020), a pioneering framework in emotion classification within textual content. GoEmotions excels in its comprehensive categorization, identifying expressions across 28 distinct emotional categories (27 emotions plus a neutral category). This framework is instrumental in enhancing sentiment analysis and interpreting a broad spectrum of human emotions, particularly in the context of concise social media posts.
In parallel, we draw insights from SODA (SOcial DiAlogues) Kim et al. (2023), a unique dataset in the realm of social dialogue systems. SODA synthesizes 1.5 million dialogues, integrating socially-grounded conversations from a pre-trained language model with contextual knowledge from the Atomic10x knowledge graph West et al. (2022). This integration addresses key challenges in dialogue agent training, namely diversity, scale, and quality, thereby surpassing traditional human-authored corpora in terms of consistency, specificity, and naturalness. SODA’s contributions are pivotal in developing dialogue agents that are more authentic and coherent.
Additionally, our work touches upon the critical concept of predictive uncertainty in Large Language Models (LLMs) Jing et al. (2024); Su et al. (2024). This uncertainty bifurcates into aleatoric uncertainty, rooted in inherent randomness in data, and epistemic uncertainty, originating from incomplete data knowledge. Recent scholarly efforts have been directed towards addressing these uncertainties in deep neural models, highlighting their significance in enhancing model reliability.
A notable development in this domain is presented in Li et al. (2023), which introduces the CUE (Contextual Uncertainty Elimination) framework. This innovative approach leverages a Variational Auto-encoder to modify latent text representations, thereby elucidating the sources of predictive uncertainty in LLM classifiers. CUE extends beyond mere identification of influential input tokens; it recalibrates text representations to provide a novel lens for interpreting uncertainty in LLM-based text classifiers. This advancement is pivotal in our understanding of LLM behavior, particularly in enhancing model transparency and developing effective strategies for uncertainty mitigation.
@@ 3 Limitations and Hypotheses
Firstly, the emotion categories in GoEmotions dataset have uneven distributions in nature. As shown in Figure 1, some emotion categories such as "admiration", "approval", and "annoyance" have more than 10k examples in the dataset, while some minority classes such as "pride", "relief", and "grief" only have less than 1k examples. As a result, there will be an approximately 10 times difference in training data sample size between some majority classes and minority classes, causing a nonnegligible difference on single category classification performance. For instance, Demszky and his team fine-tuned a BERT model on GoEmotions dataset and the "grief" category which has the least sample size in training set achieved all zeros across different evaluation metrics Demszky et al. (2020).
Figure 1: Emotion Categories Ordered by Number of Examples in the DatasetDemszky et al. (2020)
@@
    fine-tune RoBERTa model on GoEmotions dataset and compare the performance with BERT benchmark
@@
    apply three data augmentation techniques on original training dataset before fine-tuning RoBERTa, comparing the variance in performance
@@
    fine-tune RoBERTa on CARER dataset Saravia et al. (2018) first and then fine-tune the trained model on augmented GoEmotions training dataset, comparing the variance in performance
@@ 4.1 Fine-tuning BERT on GoEmotions
The first set of experiments in the original GoEmotions paper Demszky et al. (2020) is to fine-tune a pre-trained bert-base-cased model on the GoEmotions dataset with three different taxonomy hierarchies. The first hierarchy we called it the original taxonomy, which contains 27 emotion labels and a neutral label. The second hierarchy is called the grouped taxonomy. The grouped taxonomy is generated by applying a hierarchical clustering on the original 27 labels using correlation as a distance metric, resulting in 3 categories and a neutral category. The last hierarchy is called the Ekman’s taxonomy Ekman (1992), which maps the original 27 emotion labels into 6 groups with a neutral group.
Therefore, 3 taxonomies correspond to 3 different multi-label emotion classification problem. As described in the original paper, we fine-tune a bert-base-cased model on the GoEmotions dataset three times, one for each taxonomy. The model structure will be different only in the last linear layer to fit the number of classes to predict. We keep the exactly same hyperparameter configuration as the original paper: batch_size of 16, learning_rate of 5e-5 and no learning rate scheduler. We use the AdamW optimizer Loshchilov and Hutter (2017) and cross-entropy as the loss function.
@@ 4.2 Transfer Learning Experiment
The experiment aims to demonstrate the effectiveness of transfer learning in emotion understanding across various domains and taxonomies using limited labeled data in a target domain. The researchers leverage GoEmotionsDemszky et al. (2020) as the baseline dataset for training their models.
The study uses nine datasets from the Bostan and Klinger (2018) Unified Dataset, focusing on three for detailed discussion due to their diverse domains. The chosen datasets are:
@@
    ISEARR and Wallbott (1994): This dataset includes personal reports of emotional events from 3000 individuals with different cultural backgrounds. It consists of 8,000 sentences, each labeled with one of seven emotions: anger, disgust, fear, guilt, joy, sadness, and shame.
@@
    EmoIntMarrese-Taylor and Matsuo (2017): Part of the SemEval 2018 benchmark, this dataset has 7,000 tweets with crowdsourced annotations. It focuses on the intensity of four emotions: anger, joy, sadness, and fear. The researchers use a cutoff of 0.5 to obtain binary annotations.
@@
    Emotion-StimulusMohammad et al. (2018): Based on FrameNet’s emotion-directed frames, this dataset provides annotations for 2,400 sentences. Its taxonomy includes seven emotions: anger, disgust, fear, joy, sadness, shame, and surprise.
@@ 4.3 Error Analysis
However, there are still much space for improvement. Firstly, the original GoEmotions dataset is still severely imbalanced. For instance, there are more than 15k data points for the emotion category "admiration", but less than 1k data points for the emotion category "relief". As a result, we achieved 0.67 F1 score on "admiration" category while only 0.32 F1 score on the "relief" category. We suggest that some data augmentation methods can be applied on the minority classes to further improve the model performance Su et al. (2022). Moreover, our model performance improvements mainly benefit from using more training epochs. So we believe that executing thorough and proper hyperparameter tuning process can help yield better performance as well.
4.4 A Stronger Baseline?
In comparison, we purpose to fine-tune a RoBERTa model Liu et al. (2019) on the GoEmotions training set as the new baseline to replace the BERT baseline proposed in the original paper Demszky et al. (2020). Since RoBERTa optimizes the set of pre-training tasks, the pre-training configuration, and the pre-training dataset Liu et al. (2019), it is expected to outperform the performance of the BERT baseline on GoEmotions in our use case.
We keep exactly the same hyperparameter configuration for RoBERTa as how BERT was trained in the original paper: batch_size of 16, learning_rate of 5e-5 and no learning rate scheduler. We use the AdamW optimizer Loshchilov and Hutter (2017) and cross-entropy as the loss function. Same as BERT baseline, we trained RoBERTa for 4 epochs on the training set to prevent overfitting and the performance comparison with BERT baseline is shown in Table 6.
@@ 4.5 Data Augmentation
Easy Data Augmentation (EDA)
This straightforward method proposed by Wei et al. Wei and Zou (2019) has proven to be notably effective. EDA integrates various manipulations for the words present in the initial document. In our use case, for each sentence within the training data, EDA selects and executes one of the following operations with certain predefined probability:
  • Synonym Replacement: random words are replaced by random synonyms from a dictionary e.g. WordNet Miller (1995)
@@
BERT Embeddings
To address the drawback in EDA approach, a more sophisticated way to altering the original text without hurting the meaning and context of the sentence is to use BERT embeddings Devlin et al. (2019). Essentially, the word w in the original text is substituted with words forecasted by a model like BERT, considering the surrounding context of w in the original text. In our use case, a new word will be injected to random position according to contextual word embeddings or an original word will be replaced by a new word according to contextual embedding distances with predefined probability.
@@
The third method ProtAugment Dopierre et al. (2021) shifts the focus from individual words to rephrasing the entire original text. It uses BART, a model that "merges Bidirectional and Auto-Regressive Transformers" Lewis et al. (2019), for creating paraphrases. The authors of ProtAugment tuned BART specifically for paraphrase generation, employing multiple datasets for this purpose. Additionally, they employed techniques like Diverse Beam Search Vijayakumar et al. (2018) and Back Translation Mallinson et al. (2017) to help producing diverse outputs.
@@ 4.6 Transfer Learning
In this section, we want to test the impact of introducing Transfer Learning into our experiment settings. This idea is learned from the original paper of GoEmotions Demszky et al. (2020), and we use CARER Saravia et al. (2018), another Emotion Classification Task Dataset, as additional data source.
@@ 4.6.1 CARER Overview
CARER-BERT on EDA Augmented GoEmotions 49.19% 57.97% 46.38% 50.03%
@@ 4.6.2 Experiment Setting
We pick BERT Devlin et al. (2019) as our base model, which is aligned with our other experiments, and we first finetune the base model on CARER Saravia et al. (2018). Then we train our model on 4 different Datasets: Oringinal GoEmotions, EDA Augmented GoEmotions, BERT Embedding Augmented GoEmotions, PROT Augmented GoEmotions. Note that all augmented methods are only applied to under-performanced categories.
@@
We use trainer of HuggingFace Wolf et al. (2020) as main code frame and migrate our training process to GPU devices using HuggingFace Accelerate Gugger et al. (2022).

### 09 Apr 2024 v2

@@ Abstract
1 Introduction
Recently Large Language Models have gain enormous attention in the field of Natural Language Processing. The task of detecting subtle emotions in text is a complex issue in Natural Language Processing with numerous practical uses. A major hurdle is the lack of sufficiently large and annotated datasets for emotion detection. Demszky Demszky et al., (2020) created the GoEmotions dataset, which is the largest manually annotated dataset containing 58k English Reddit comments, labeled for 27 emotion categories and Neutral. The researchers also fine-tuned a BERT model on GoEmotions as a strong baseline in transfer learning across domains and taxonomies to show the potentials of this dataset.
In the initial experiments, our team dived deep into the original paper and reproduced the results of both fine-tuning BERT model on GoEmotions dataset and applying transfer learning technique using fine-tuned BERT to other emotion classification dataset with different taxonomies. We were able to achieve 0.49 macro-average F1 score on the 28-label classification task, which was slightly improved over the best result 0.46 reported in the original paper Demszky et al., (2020). However, we believe that there is still much room for improvement in term of classification performance on GoEmotions dataset. Therefore, our main objective in this study is to utilize different methods to enhance the classification performance on the fine-grained emotion detection dataset, GoEmotions, as much as possible.
@@ 2 Related Work
Our study builds upon the foundation laid by GoEmotions Demszky et al., (2020), a pioneering framework in emotion classification within textual content. GoEmotions excels in its comprehensive categorization, identifying expressions across 28 distinct emotional categories (27 emotions plus a neutral category). This framework is instrumental in enhancing sentiment analysis and interpreting a broad spectrum of human emotions, particularly in the context of concise social media posts.
In parallel, we draw insights from SODA (SOcial DiAlogues) Kim et al., (2023), a unique dataset in the realm of social dialogue systems. SODA synthesizes 1.5 million dialogues, integrating socially-grounded conversations from a pre-trained language model with contextual knowledge from the Atomic10x knowledge graph West et al., (2022). This integration addresses key challenges in dialogue agent training, namely diversity, scale, and quality, thereby surpassing traditional human-authored corpora in terms of consistency, specificity, and naturalness. SODA’s contributions are pivotal in developing dialogue agents that are more authentic and coherent.
Additionally, our work touches upon the critical concept of predictive uncertainty in Large Language Models (LLMs) Jing et al., (2024); Su et al., (2024). This uncertainty bifurcates into aleatoric uncertainty, rooted in inherent randomness in data, and epistemic uncertainty, originating from incomplete data knowledge. Recent scholarly efforts have been directed towards addressing these uncertainties in deep neural models, highlighting their significance in enhancing model reliability.
A notable development in this domain is presented in Li et al., (2023), which introduces the CUE (Contextual Uncertainty Elimination) framework. This innovative approach leverages a Variational Auto-encoder to modify latent text representations, thereby elucidating the sources of predictive uncertainty in LLM classifiers. CUE extends beyond mere identification of influential input tokens; it recalibrates text representations to provide a novel lens for interpreting uncertainty in LLM-based text classifiers. This advancement is pivotal in our understanding of LLM behavior, particularly in enhancing model transparency and developing effective strategies for uncertainty mitigation.
@@ 3 Limitations and Hypotheses
Firstly, the emotion categories in GoEmotions dataset have uneven distributions in nature. As shown in Figure 1, some emotion categories such as "admiration", "approval", and "annoyance" have more than 10k examples in the dataset, while some minority classes such as "pride", "relief", and "grief" only have less than 1k examples. As a result, there will be an approximately 10 times difference in training data sample size between some majority classes and minority classes, causing a nonnegligible difference on single category classification performance. For instance, Demszky and his team fine-tuned a BERT model on GoEmotions dataset and the "grief" category which has the least sample size in training set achieved all zeros across different evaluation metrics Demszky et al., (2020).
Figure 1: Emotion Categories Ordered by Number of Examples in the DatasetDemszky et al., (2020)
@@
The third limitation results from the reliance on traditional language models, such as BERT, in the GoEmotions paper which was published four years ago. There is a lack of exploration into how modern large language models (LLMs) like GPT-4 and Llama would perform on emotion detection datasets and into their potential for understanding and interpreting human emotions.
@@
  4. 4.
    Modern LLMs like GPT-4 and Llama will, under zero-shot settings, outperform traditionally fine-tuned Language Models. This paper is the first to evaluate the capabilities of LLMs on the GoEmotions dataset.
@@
    Reproduce the experiments from GoEmotions Dataset paper Demszky et al., (2020) and compare the performance with the published benchmark. This step is to ensure our experiment environment are correctly set because it has been 4 years and the original paper trained the networks using TensorFlow framework whereas we are using HuggingFace and PyTorch.
@@
    Fine-tune RoBERTa model on GoEmotions dataset and compare the performance with BERT benchmark.
@@
    Apply three data augmentation techniques on original training dataset before fine-tuning RoBERTa, comparing the variance in performance.
  4. 4.
    Fine-tune RoBERTa on CARER dataset Saravia et al., (2018) first and then fine-tune the trained model on augmented GoEmotions training dataset, comparing the variance in performance.
  5. 5.
    Introduce modern LLMs like Llama Touvron et al., (2023) and GPT-4 OpenAI and et al, (2024) and test them on a subset of GoEmotions and compare them with BERT benchmark. This paper is the first to put LLMs in test of GoEmotions dataset.
@@ 4.1 Fine-tuning BERT on GoEmotions
The first set of experiments in the original GoEmotions paper Demszky et al., (2020) is to fine-tune a pre-trained bert-base-cased model on the GoEmotions dataset with three different taxonomy hierarchies. The first hierarchy we called it the original taxonomy, which contains 27 emotion labels and a neutral label. The second hierarchy is called the grouped taxonomy. The grouped taxonomy is generated by applying a hierarchical clustering on the original 27 labels using correlation as a distance metric, resulting in 3 categories and a neutral category. The last hierarchy is called the Ekman’s taxonomy Ekman, (1992), which maps the original 27 emotion labels into 6 groups with a neutral group.
Therefore, 3 taxonomies correspond to 3 different multi-label emotion classification problem. As described in the original paper, we fine-tune a bert-base-cased model on the GoEmotions dataset three times, one for each taxonomy. The model structure will be different only in the last linear layer to fit the number of classes to predict. We keep the exactly same hyperparameter configuration as the original paper: batch_size of 16, learning_rate of 5e-5 and no learning rate scheduler. We use the AdamW optimizer Loshchilov and Hutter, (2017) and cross-entropy as the loss function.
@@ 4.2 Transfer Learning Experiment
The experiment aims to demonstrate the effectiveness of transfer learning in emotion understanding across various domains and taxonomies using limited labeled data in a target domain. The researchers leverage GoEmotionsDemszky et al., (2020) as the baseline dataset for training their models.
The study uses nine datasets from the Bostan and Klinger, (2018) Unified Dataset, focusing on three for detailed discussion due to their diverse domains. The chosen datasets are:
@@
    ISEARR and Wallbott, (1994): This dataset includes personal reports of emotional events from 3000 individuals with different cultural backgrounds. It consists of 8,000 sentences, each labeled with one of seven emotions: anger, disgust, fear, guilt, joy, sadness, and shame.
@@
    EmoIntMarrese-Taylor and Matsuo, (2017): Part of the SemEval 2018 benchmark, this dataset has 7,000 tweets with crowdsourced annotations. It focuses on the intensity of four emotions: anger, joy, sadness, and fear. The researchers use a cutoff of 0.5 to obtain binary annotations.
@@
    Emotion-StimulusMohammad et al., (2018): Based on FrameNet’s emotion-directed frames, this dataset provides annotations for 2,400 sentences. Its taxonomy includes seven emotions: anger, disgust, fear, joy, sadness, shame, and surprise.
@@ 4.3 Analysis and Insights
However, there are still much space for improvement. Firstly, the original GoEmotions dataset is still severely imbalanced. For instance, there are more than 15k data points for the emotion category "admiration", but less than 1k data points for the emotion category "relief". As a result, we achieved 0.67 F1 score on "admiration" category while only 0.32 F1 score on the "relief" category. We suggest that some data augmentation methods can be applied on the minority classes to further improve the model performance Su et al., (2022). Moreover, our model performance improvements mainly benefit from using more training epochs. So we believe that executing thorough and proper hyperparameter tuning process can help yield better performance as well.
4.4 Is Fine-tuned RoBERTa A Stronger Baseline?
In comparison, we propose to fine-tune a RoBERTa model Liu et al., (2019) on the GoEmotions training set as the new baseline to replace the BERT baseline proposed in the original paper Demszky et al., (2020). Since RoBERTa optimizes the set of pre-training tasks, the pre-training configuration, and the pre-training dataset Liu et al., (2019), it is expected to outperform the performance of the BERT baseline on GoEmotions in our use case.
We keep exactly the same hyperparameter configuration for RoBERTa as how BERT was trained in the original paper: batch_size of 16, learning_rate of 5e-5 and no learning rate scheduler. We use the AdamW optimizer Loshchilov and Hutter, (2017) and cross-entropy as the loss function. Same as BERT baseline, we trained RoBERTa for 4 epochs on the training set to prevent overfitting and the performance comparison with BERT baseline is shown in Table 6.
@@ 4.5 Data Augmentation
Duplication Data Augmentation (DDA)
This straightforward method proposed by Wei et al. Wei and Zou, (2019) has proven to be notably effective. DDA integrates various manipulations for the words present in the initial document. In our use case, for each sentence within the training data, DDA selects and executes one of the following operations with certain predefined probability:
  • Synonym Replacement: random words are replaced by random synonyms from a dictionary e.g. WordNet Miller, (1995)
@@
Traditional Language Model Embeddings
To address the drawback in DDA approach, a more sophisticated way to altering the original text without hurting the meaning and context of the sentence is to use BERT embeddings Devlin et al., (2019). Essentially, the word w in the original text is substituted with words forecasted by a traditional Language Model (LM) like BERT, considering the surrounding context of w in the original text. In our use case, a new word will be injected to random position according to contextual word embeddings or an original word will be replaced by a new word according to contextual embedding distances with predefined probability.
@@
The third method ProtAugment from Dopierre et al., (2021) shifts the focus from individual words to rephrasing the entire original text. It uses BART, a model that "merges Bidirectional and Auto-Regressive Transformers" Lewis et al., (2019), for creating paraphrases. The authors of ProtAugment tuned BART specifically for paraphrase generation, employing multiple datasets for this purpose. Additionally, they employed techniques like Diverse Beam Search Vijayakumar et al., (2018) and Back Translation Mallinson et al., (2017) to help producing diverse outputs.
@@ 4.6 Transfer Learning
In this section, we want to test the impact of introducing Transfer Learning into our experiment settings. This idea is learned from the original paper of GoEmotions Demszky et al., (2020), and we use CARER Saravia et al., (2018), another Emotion Classification Task Dataset, as additional data source.
@@ 4.6.1 CARER Overview
CARER-BERT on DDA Augmented GoEmotions 49.19% 57.97% 46.38% 50.03%
@@ 4.6.2 Experiment Setting
We pick BERT Devlin et al., (2019) as our base model, which is aligned with our other experiments, and we first finetune the base model on CARER Saravia et al., (2018). Then we train our model on 4 different Datasets: Oringinal GoEmotions, DDA Augmented GoEmotions, BERT Embedding Augmented GoEmotions, PROT Augmented GoEmotions. Note that all augmented methods are only applied to under-performanced categories.
@@
We use trainer of HuggingFace Wolf et al., (2020) as main code frame and migrate our training process to GPU devices using HuggingFace Accelerate Gugger et al., (2022).
@@ 4.6.3 Transfer Learning Results
4.7 Modern Large Language Models on GoEmotions
As discussed above as the fifth hypothesis, we aim to test the performance of the modern LLMs on Emotione Detection task, where m̈odern LLMsr̈efers to LLMs that have billion-level parameter size et al, (2023), differing to those with million-level like BERT tested in the original GoEmotions paper. The experiment settings for testing modern LLMs is that we use a subset of the GoEmotions dataset, 1000 data points, as input to LLMs with the following prompts:
"I want see your performance in emotion detection task. The setting aligns with the GoEmotions paper with title of "GoEmotions: A Dataset for Fine-Grained Emotion Classification". There are in total 28 categories as emotion labels and every data entry can have one or more of the 28 categories. I will give you a number of sentences that may or may not contain certain emotions inside, and your job is to label the input sentences with the given 28 categories like the setting of GoEmotions paper. Please organize the labels into a string of words with comma as separator. The output should be a csv table with 3 columns where every row contains the id, the given sentence, and your label."
We use the above text as prompt and we call OpenAI’s API for ChatGPT-4 using 30 sentences as a batch input in every call. We then collect their response and calculate the accuracy. The calculation on metrics is different than what we did to BERT and RoBERTa and we will address this in the Analysis section. Note that we particularly want to test their zero-shot performance on emotion detection task thus there is no training input. We use GPT-4 as the modern LLM and we skip including Llama in the scope as it cannot tell the 28 categories correctly and it can only tell emotions from positive, negative, and neutral.
@@ 5.2 Transfer Learning Analysis
5.3 Modern LLMs on GoEmotions Analysis
As shown in the table 12, in zero-shot setting, comparing to fine-tuned BERT and RoBERTa, the performance of GPT-4 is much lower in accuracy, precision, recall and F1-score. Note that here we keep the same way of calculating metrics of using macro average on every emotion class. We summarize the reasons of such low performance as followings:
Hallucination
. The problem of hallucination is well-known to be associated with generative models and LLMs Jing et al., (2024), and we found in our case, GPT-4 will classify a sentence with a emotion label that is not included in the 28 categories. Here is an example: the given sentence is "10 life hacks you didn’t know you needed in your life." and GPT-4 put "curiosity" and "informative" as emotion labels to it, where informative is not one of the designated categories. Similar problems happen to quite a few data points. In fact, to our test of 1000 data points of GoEmotions dataset, there are in total 89 examples with emotion labels that are made up by GPT-4 and not included in the 28 categories in reality. This is to say, by using 30 data points as batch input, there are almost one made-up label in every 10 given input. But we have emphasize that 28-class classification task is much harder than 3 classes of positive, negative, and neutral as labels in other common emotion detection task.
Over-labelling
. We found that almost every predicted sentence have at least 2 labels in GPT-4’s response whereas multi-label data points are much less common in the original dataset of GoEmotions. We calculate the number of the examples that have more predicted labels than the true label, and we found up to 812 out of 1000 data points. The majority of such examples exhibit the problem of correctly predicting the true labels while adding a new emotion label that is similar or close in semantic meaning of the predicted ones. This would decrease the result in precision, recall, and F1 score.
Over-interpretation
. Often the case is GPT-4 will give a response that overly interprete the meaning of the sentence. For example, the given sentence is "Ask him out for a drink.", and in normal occasion this is simply a proposal to ask somebody out. The true label for this sentence in the dataset is "neutral" whereas the response from GPT-4 is "desire" and "optimism". This is clearly over-interpretation of the meaning of the sentence.
In summary, we are the first to apply modern LLMs to GoEmotions dataset, and we observed 3 challenges when employing GPT-4 for emotion classification tasks, especially when benchmarked against fine-tuned models like BERT and RoBERTa.